# Decision-dependent Robust Charging Infrastructure Planning

### v1

@@ 4 Fix-and-optimize heuristic based on high charging demand forecasts
   Solve the proposed optimization problem with deterministic $pp_{i,d,t}$ as the initial input (master problem); $\triangleright$ Warm start if the solution of the master problem available;
   Obtain scheduling outcomes $y_{i,j,d,t}$ and charging profiles $p^{ch}_{i,j,d,t}$;
  Sample parking durations $\hat{pp}_{i,d,t}$ from the uncertainty set $\Phi(\mu,\epsilon)$;
 for $i\in\mathcal{I},\;t\in\mathcal{T},\;d\in\mathcal{D},\;j\in\mathcal{J}$ do
      Compute the new upper bound of charging power:
$p^{ch^{\prime}}_{i,j,d,t}\leftarrow\sum_{j\in\mathcal{J}}\Delta t\cdot P^{j}\cdot y_{i,j,d,t}\cdot\hat{pp}_{i,d,t}\cdot\eta^{j}\;$
    if $\big(p^{ch}_{i,0,d,t}\geq p^{ch^{\prime}}_{i,0,d,t}+\kappa^{v}\big)\land(\forall i,d,t\in\overline{\mathcal{S}^{Z}})$ or $p^{ch}_{i,1,d,t}\geq p^{ch^{\prime}}_{i,1,d,t}$ then
@@
Post-process the solution by removing fast charging results in zones where the number of fast chargers exceeds two;
Re-solve the optimization problem using the sampled values $\hat{pp}_{i,d,t}$ and incorporating the charger installation constraints derived from the processed solution;
Algorithm 1 Fix-and-optimize heuristic under effective parking duration uncertainty
This iterative process continues until the maximum number of iterations is reached. Afterward, we post-processed the optimization results by removing zones with more than two fast chargers. The remaining fast charger installations were treated as hard constraints, while the solutions for slow chargers were used as lower bounds. Ultimately, the installation plan for all zones and charger types was determined.
@@ 5 Application
We applied our model to an open-pit mining site with 203 light-duty trucks in operation, which are planned to be replaced by battery EVs. Seven special-use trucks, such as ambulances, that will not be electrified, along with another seven trucks dispatched fewer than 10 days, are excluded from the study. The remaining 189 trucks will be replaced with electric trucks equipped with 100 kWh batteries. We consider the following two types of chargers, as listed in Table 1. The capital cost of fast chargers includes the installation cost. We set the penalty of low SoC and charging time to be one in the objective function. The charging efficiency could vary with the charging type, outdoor temperature, and charging power \@BBOPcite\@BAP\@BBNSevdari et al. (2023)\@BBCP. This paper uses a fixed value for charging efficiency. Overall, fast charging tends to be more efficient than slow charging, as it uses a high-efficiency rectifier located outside the truck \@BBOPcite\@BAP\@BBNArena et al. (2024)\@BBCP. We therefore set the charging efficiency of slow and fast charging to be 90% and 95%.
@@
Capital costs (\$/charger) 1,500 \@BBOPcite\@BAP\@BBNFord (2025)\@BBCP 38,000 \@BBOPcite\@BAP\@BBNPrimecomtech (n.d.)\@BBCP
@@
The charging planning process begins by identifying potential charging sites based on the stop frequency of all trucks (Section 5.1). Next, the parking window, effective parking duration, and energy consumption for each truck are extracted (Section 5.2). Using this information, the DS and IPS methods are applied to plan across the full-year horizon. If the RO formulation is used, a decision-dependent uncertainty set is constructed from the historical samples (Section 5.3). High-consumption vehicles with infeasible trips are then filtered out (Section 5.4). Finally, chargers are planned through the proposed model in the deterministic or stochastic formulation (Section 5.5).
@@ 5.1 Identifying potential EV charging points
We used GPS records of 189 light-duty trucks at the mining site over one year. Their GPS data records include coordinates (i.e., longitude and latitude) and speed at each time step. We then grid-divided the entire mining site into 100-meter $\times$ 100 100-meter square regions and counted the stop frequency in each square. We assumed that a truck stops if it moves less than 50 meters within a square region for at least five minutes. Comparing the number of stop frequencies in each square, we identified eight EV charging points with the highest number of stops throughout the year. Fig. 5 shows eight EV charging points identified, and the dashed gray lines indicate the roads connecting them. Fig. 6 shows the total stopping time of the entire fleet ($\times$ 1,000 hours), including both short parking durations (<< 4 hours) and long parking durations ($\geq$ 4 hours), along with the annual number of trucks stopping at each zone. The detailed statistics of eight charging zones are listed in Table 6 in Appendix 9.1. The zone 1 (roadhouse) and zone 4 (equipment maintenance) have a large part of long duration time greater than 4 hours, which means that most of the overnight parking happens at zones 1 and 4.
@@ 5.2 Summary of input datasets
5.3 Samples for constructing robust uncertainty sets
To construct a robust uncertainty set for heterogeneous effective parking durations, we first collected all parking durations and parking window data across the year. Fig. 8 (a) illustrates the total time of actual parking duration versus the the total time of the full parking window per truck per day at eight zones. In each zone, a portion of the parking window cannot be used for charging, for example, when the truck is moving to find a parking spot. Fig. 8 (b) presents the distribution of parking durations across eight zones, showing that zones 6, 7, and 8 exhibit greater variability compared to the others.
Figure 8: (a) Actual parking duration versus full parking windows duration per day per truck and (b) box plots of actual parking duration distributions of eight zones
Figure 9: (a) Mean and (b) variance of effective parking durations in minutes of 189 trucks by hour of the day in the charging zone 1
We found that the actual parking duration at the mining site is strongly related to trucks, hour of the day, and charging zones, but has almost no correlation with day of the week or day of the month. This is because the mining site operates continuously with the same pattern throughout the year. We then calculate the mean and standard deviation of parking durations pertaining to trucks, hours of the day, and zones. Fig. 9 shows the mean and variance of effective parking duration in minutes by hour of the day in the charging zone 1. Most trucks operating around the mining site have zero means and variances in their actual parking duration during the day.
@@ 5.4 Set up
We prepared three inputs based on the dataset of the year 2023, for the DS and IPS approaches, respectively. For the DS approach, we defined 10 statistical criteria, including total fleet distance, number of active trucks, and total fleet energy consumption. Based on these criteria, we ranked all 365 days in a year and selected the top-ranking days. We experimented with different numbers of selected days, ranging from 10 to 40, and observed that charger installation results stabilized once the selection reached 30 days. Accordingly, we chose 39 days, including days adjacent to the top 10 ranked days, as some vehicles charge every two or three days, and representative sub-periods need to extend beyond battery cycles. The detailed selection criteria and procedures are provided in Appendix 9.2. For the IPS approach, we divided the full-year dataset into monthly subsets, incorporating overlapping days between consecutive months to maintain trip continuity. This procedure produced 12 monthly datasets, each spanning 34 days.
We conduct a preliminary simulation before optimization, in which each EV uses straightforward fast charging throughout the optimization horizon. In each simulation, a truck starts with a random SoC and is then fully charged at each parking site. At the end of the optimization horizon, we track whether they could restore their original SoC. If the original SoC cannot be recovered, the truck is identified as a high consumption vehicle (HCV), which cannot be feasibly electrified if it sticks to the current operational scheduling. This also includes vehicles not passing the eight charging zones. The optimization is subsequently performed exclusively on the rest of the trucks, and we recorded the number of HCVs excluded from the optimization model in each simulation. All experiments were solved using the Gurobi commercial solver. We set the accepted optimal gap for this MILP to be 1%, less than the capital cost of one slow charger. All problems were solved on a computer equipped with an 8-core CPU running at 3.4 GHz and 160 GB of RAM.
@@ 5.5.1 The impacts of abandonment, range anxiety, and effective parking duration
We first evaluate four deterministic cases to analyze the impacts of charging abandonment, range anxiety, and parking duration restrictions. Using the DS method, we draw on 39 selected representative days for the input datasets. The total energy consumption of the entire vehicle fleet for the 39-day optimization horizon is 83,050 kWh, which is around 11.33 kWh per day per truck. The total parking time of the entire vehicle fleet is 124,639 hours, which is around 17 hours per day per truck.
  • Case 1 - Benchmark: Trucks are prohibited from waiting in any zone under normal circumstances. However, when a vehicle SoC drops to below 30%, it is permitted to pause for 30 minutes to recharge. Two designated zones (i.e., zone 1 and zone 4) accommodate the overnight charging with unlimited waiting time. The model incorporates the specific parking duration for each parking window (i.e., $\frac{5}{30}\leq pp\leq 1$).
  • Case 2 - Full parking time: This case maintains the same conditions as the benchmark case, except that it has no specific parking duration constraints for each parking window (i.e., $pp=1$).
@@
  • Case 4 - No range anxiety: This case maintains the same conditions as the benchmark case, except that it has no related constraints or penalty related to the range anxiety, in order words, there is no penalty for low SoC (i.e., $P^{low}=1$). Trucks are prohibited from waiting in any zone under normal circumstances, regardless of their SoC. In other words, the maximum waiting time is set to zero (i.e., $T^{max}_{i,d,t}=0$).
Table 2 summarizes optimization results in four deterministic cases under different assumptions about abandonment, range anxiety, and effective parking duration. The benchmark case (case 1) has a mixed type of chargers, which are 72 slow chargers and 7 fast chargers. Compared to the benchmark case, case 2 shows that a less number of fast chargers are needed if the model does not consider effective charging time in each parking time slot. In case 3, an additional number of 20 slow chargers are installed if there is no special zone allowing for overnight charging. In case 4, if the model does not consider range anxiety, seven additional slow chargers need to be installed to accommodate more frequent abandonment without waiting time.
Instances
# of
HCVs
# of slow /
fast chargers
Installation
cost (\$)
Scheduling
penalty (\$)
Gap (%) Time (h)
Case 1 1 72/6 336,000 38,468 0.084 2.98
Case 2 1 74/3 187,000 39,622 0.116 57.32
Case 3 1 92/6 366,000 85,200 0.042 15.77
Case 4 1 79/6 346,500 38,077 0.079 2.94
Table 2: Summary of optimization results in four deterministic cases: benchmark, no parking time restriction, no overnight charging, and no range anxiety
Figure 10: Zone-wise differences in the number of slow and fast chargers of cases 2-4 compared with the benchmark
Fig. 10 illustrates the differences in the numbers of fast and slow chargers in cases 2–4, compared with the benchmark. Across four cases, the number of installed chargers is proportional to the total duration of vehicle stops at each charging zone, as presented in Fig. 8 (a), and zones 4 and 5 have the highest number of chargers. In case 1, setting the special zones does not relax the requirement of the fast charger installation in these zones, as in zones 1 and 4. In case 3, without considering the overnight charging, zones 1 and 4 will have a high number of slow chargers. In case 4, without considering the range anxiety, zones 5-7 will have a higher number of slow chargers, meaning vehicles passing through these three zones have a relatively low SoC than other zones.
Instances
Total slow
charging
hours
Total fast
charging
hours
Total
abandonment
hours
Avg. charging
hours
(truck per day)
Avg. charging
power
Case 1 3677.0 1352.5 26371.0 0.69 16.51
Case 2 5779.0 555.5 26430.5 0.86 13.11
Case 3 3493.5 1418.5 69353.0 0.67 16.90
Case 4 3116.5 1248.5 26947.0 0.60 19.02
Table 3: Summary of the scheduled charging hours, abandonment hours, and charging power across four deterministic cases
Figure 11: The total number of (a) charging hours and (b) abandonment hours of the entire truck fleet in a day across four instances
Figure 12: The total number of (a) charging hours and (b) abandonment hours of the truck fleet in different hours of a day in four deterministic cases
@@
We calculate the charging power, charging hours, and abandonment hours across four deterministic cases, summarized in Table 3. The detailed scheduling solution of four cases are presented in Appendix 10. Case 2, which assumes full parking duration, results in the highest total charging hours but the lowest average charging power due to having the fewest fast chargers. Case 3, which excludes overnight charging, produces the greatest number of abandonment hours. Case 4, which removes range anxiety, leads to the fewest charging hours and the highest average charging power.
Figs. 11 (a) and (b) show the zone-wise comparison of charging and abandonment time, respectively. In case 2, with the full parking duration, zones 6 and 8 have a higher charging time. In case 3, without the overnight charging, zones 1 and 4 have a significantly high number of abandonment hours. Figs. 12 (a) and (b) show the total number of charging and abandonment hours of the total truck fleet in different hours of a day in four cases, respectively. Across four instances, the lowest number of charging hours appears at 12:00, and the highest number of charging hours appears at 22:00. If the model does not assign two special zones for overnight charging, the number of abandonment hours will increase by more than a factor of 2.
@@
Fig. 13 (a) illustrates the average SoC of all trucks over 39 days, ranked from the highest to lowest, and Fig. 13 (b) presents the overall battery SoC (maximum, minimum, and average) of the entire fleet in four cases. Among four cases, case 4 (no range anxiety) yields the lowest average SoC ($28.4\%$), whereas case 2 with full parking duration achieves the highest average SoC ($45.3\%$). When the effective charging time constraint is applied, in other words, comparing case 2 to case 1 benchmark, part of the fleet experiences a sharp drop in the average SoC to very low levels ($\leq 20\%$), due to the limited charging duration in the parking windows.
5.5.2 Robust infrastructure planning based on the whole year dataset
[…]
To ensure robust infrastructure planning throughout the year, we optimize the charging infrastructure using two solving approaches: the DS and IPS methods, as described in Section 3. For the DS method, we selected a 39-day horizon and imposed a maximum time limit of 40,000 seconds to solve each instance, recording the gap whenever the solver failed to converge within this limit. For the IPS method, we set a 6-hour time limit for each monthly instance. In addition, we applied the proposed fix-and-optimize heuristic to the 39-day instance under the DS framework, with the initial master problem aligned with the deterministic benchmark case.
Instances
($\sigma$, $\gamma$)
Solving
methods
# of
HCVs
# of slow/fast
chargers
Installation
costs (\$)
Scheduling
penalty (\$)
Gap (%)
Time (h)
(0, 1) DS 19 68/6 330,000 35,678 9.06 11.1
DS - h 19 68/6 330,000 35,679 0.07 2.40
IPS 48 77/5 305,500 33,466 0.09 11.3
(1, 1) DS 19 68/6 330,000 35,658 9.04 11.1
DS - h 19 69/6 331,500 36,065 0.05 2.15
IPS 48 75/7 378,500 33,331 0.05 12.0
(2, 1) DS 19 68/6 330,000 35,736 9.02 11.1
DS - h 19 69/6 331,500 35,863 0.05 5.20
IPS 48 76/8 418,000 33,402 0.05 23.8
(2, 0.5) DS 19 69/5 293,500 35,931 0.06 7.42
DS - h 19 68/6 330,000 35,715 0.06 3.29
IPS 48 74/8 415,000 22,541 0.03 29.8
Table 4: Summary of optimization results across four instances using DS, DS with heuristic (DS-h), and IPS methods. For the IPS method, the reported gap is the optimal gap of the final instance in the iterative planning
Figure 14: The total number of chargers in each month using the IPS method in four instances (in colored lines) versus the total number of trucks parking and charged (in gray bars) in all eight zones
[…]

### v2

@@ 3.2 Fix-and-optimize heuristic based on high charging demand forecasts
   Solve the master problem $\mathcal{M}$; $\triangleright$ Warm start at the first step if the solution available;
   Obtain scheduling outcomes $\mathbf{y}=(y^{j}_{i,d,t})_{(i,d,t)\in\mathcal{S},j\in\mathcal{J}}$ and charging power $\mathbf{p}^{ch}=(p^{ch})_{(i,d,t)\in\mathcal{S}}$;
 for $(i,d,t)\in c\;,j\in\mathcal{J}$ do
      Compute the lower bound of parking durations $\hat{pp}_{i,d,t}=\mu_{c}-\sigma_{c}$ for the cluster $c$;
    $\triangleright$ Applying scaling factors $k^{j}$ if RO-DDU ;
      Compute the updated upper bound of charging power $p^{ch^{\prime}}_{i,j,d,t}$
$p^{ch^{\prime}}_{i,j,d,t}\leftarrow\sum_{j\in\mathcal{J}}\Delta t\cdot P^{j}\cdot y^{j}_{i,d,t}\cdot\hat{pp}_{i,d,t}\cdot\eta^{j}\;$
    if Violations are detected then
@@
Post-process the solution by discarding part of the solution and re-solve the problem with $\hat{pp}_{i,d,t}$;
Algorithm 1 Fix-and-optimize heuristic under parking duration uncertainty
As illustrated in Algorithm 1, we first initialize the master problem with the starting values $pp_{i,d,t}$. Next, we compute the lower bound of parking duration $\hat{pp}_{i,d,t}$ based on the cluster $c$ to replace the initial values and compute the updated bounds of the charging power. We then compare the updated upper bounds to the required charging power in the master problem $\mathcal{M}$ to identify any violations. A violation occurs if the required charging power exceeds the upper bound by more than a threshold $\kappa^{v}$ and the parking window does not belong to a special zone. The threshold $\kappa^{v}$ is initialized at 2 and decays exponentially across iterations. When a violation is detected, a fast-charging scheduling constraint $y_{i,1,d,t}=1$ is added for that parking window. These constraints are subsequently incorporated into the master problem, after which the model is resolved from the first step.
This iterative process continues until the maximum number of iterations is reached. Afterward, we post-processed the optimization results by discarding part of the solution, for instance, a high number of fast chargers in one zone. The remaining solutions for fast and slow chargers are then used as lower bounds to resolve the model. Ultimately, the installation plan for all zones and charger types was determined.
@@ 4 Application
We applied our model to an open-pit mining site with 203 light-duty trucks in operation, planned for replacement with BEVs. Seven special-use trucks, such as ambulances, that will not be electrified, along with another seven trucks dispatched fewer than 10 days, are excluded from the study. The remaining 189 trucks will be replaced with electric trucks equipped with 100 kWh batteries. We consider the following two types of BEV chargers, as listed in Table 2. The capital cost of fast chargers does not include installation costs, as installation is location-dependent. We set the penalties for low SoC and charging time to be equal in the objective function. Although the charging efficiency could vary with the charging type, outdoor temperature, and charging power \@BBOPcite\@BAP\@BBNSevdari et al. (2023)\@BBCP, this paper uses a fixed value for charging efficiency. Overall, fast charging is more efficient than slow charging because it uses a high-efficiency rectifier located outside the truck \@BBOPcite\@BAP\@BBNArena et al. (2024)\@BBCP. We therefore set the charging efficiencies for slow and fast charging to 90% and 95%.
@@
Capital costs (\$/charger) 1,500 \@BBOPcite\@BAP\@BBNFord (2025)\@BBCP 38,000 \@BBOPcite\@BAP\@BBNPrimecomtech (2025)\@BBCP
@@
The charging planning process begins by identifying potential charging sites based on the stop frequency of all trucks (Section 4.1). Next, the parking window, effective parking duration, and energy consumption of each truck are extracted (Section 4.2). Then, hierarchical clustering based on the rearranged time-zone features is applied to the parking-duration samples to construct feature-based uncertainty sets (Section 4.3). High-consumption vehicles (HCVs) with infeasible trips are then filtered out (Section 4.4). Finally, the proposed model plans chargers using different scheduling operations (Section 4.5).
@@ 4.1 Identifying potential EV charging points
We used GPS data from 189 light-duty trucks at a mining site over one year. Their GPS data records include coordinates (i.e., longitude and latitude) and speed at each time step. We then grid-divided the entire mining site into 100-meter $\times$ 100-meter square regions and counted the stop frequency in each square. We assumed that a truck stops if it moves less than 50 meters within a square region for at least five minutes. By comparing the number of stop frequencies per square, we identified eight EV charging points with the highest annual stop frequency. Fig. 5 shows eight EV charging points identified, and the dashed gray lines indicate the roads connecting them. Fig. 6 shows the total stopping time of the entire fleet ($\times$ 1,000 hours), including both short parking durations (<< 4 hours) and long parking durations ($\geq$ 4 hours), along with the annual number of trucks stopping at each zone. The detailed statistics of eight charging zones are listed in Table 10 in Appendix 8.1. Zones 1 (roadhouse) and 4 (equipment maintenance) account for a large share of long-duration time greater than 4 hours, indicating that most overnight parking occurs in zones 1 and 4.
@@ 4.2 Summary of input datasets
4.3 ML clustering for constructing feature-based robust sets
Figure 8: Parking duration distributions in a year in terms of four features: (a) truck category, (b) half-hour of a day, (c) zones, and (d) day of the month
Figure 9: Average parking duration heatmap for the selected 20 trucks, constructed using rearranged zone–hour features. The heatmap displays only the first 10 hours for each of the eight zones, with zone boundaries indicated by black dashed lines. The hierarchical clustering dendrogram, based on feature distances, is shown on the left.
To construct a feature-based, robust uncertainty set for heterogeneous parking duration, as illustrated in Section 2.7.1, we first collected all parking durations in a year. Fig. 8 (a) - (d) illustrate parking duration distributions in terms of four features, truck category, half-hour of the day, zone, and day of the month. We observed correlations between truck categories, half-hour of the day, and zones and parking duration. We then rearranged the time-zone features based on their combinations, and this gives a total number of $8\text{zones}\times 48\text{time slots}=384$ features.
We then apply the proposed hierarchy clustering to group trucks based on the re-arranged time-zone features, as demonstrated in Section 2.7.1. Before clustering, we normalized the parking duration of these trucks using z-scores. The z-score standardizes the data for each truck to a mean of zero with a standard deviation of one. In other words, the average parking duration of a truck is normalized to one, and the duration that is higher or lower than the average value is a positive or negative value, respectively. Fig. 9 shows the normalized parking duration heatmap for the selected 20 trucks, constructed using rearranged zone–hour features, with the hierarchical clustering dendrogram on the left. We first grouped trucks by category. Using the category information, hierarchical clustering yields three subgroups per category, for a total of 24 subgroups. We also compare these methods with the other three: hierarchical clustering without category information, k-means, and grouping each truck separately. These comparisons and clustering results are detailed in Appendix 9.
@@ 4.4 Set up
We prepared three inputs for the selective and full-year datasets for 2023. For the selective day approach, we defined 10 statistical criteria to select 39 days, including total fleet distance, number of active trucks, and total fleet energy consumption. Based on these criteria, we ranked all 365 days in a year and selected the top-ranking days. We tested different numbers of selected days, from 10 to 40, and observed that charger installation results stabilized at 30 days. Accordingly, we selected 39 days, including days adjacent to the 10 top-ranked days, because some vehicles charge every 2 or 3 days, and representative sub-periods must extend beyond battery cycles. The detailed selection criteria and procedures are provided in Appendix 8.2.2. For the full year using rolling-horizon scheduling, we divided the full-year dataset into monthly subsets, with overlapping days between consecutive months to maintain trip continuity. This produces 12 monthly sub-datasets, each spanning 34 days.
We conduct a preliminary simulation before optimization, in which each EV uses straightforward fast charging throughout the optimization horizon. In each simulation, a truck starts with a random SoC, and is then fully charged at each parking site. At the end of the optimization horizon, we track whether they could restore their original SoC. If the original SoC cannot be recovered, the truck is identified as a HCV, which is not feasible to be electrified if they stick to the current operational scheduling. This also includes vehicles that do not pass eight charging zones. The optimization is subsequently performed exclusively on the remaining trucks, and we recorded the number of HCVs excluded from the optimization model for each simulation. All experiments were solved using the Gurobi commercial solver. We set the accepted optimal gap for this MILP to be 1%, less than the capital cost of one slow charger. All problems were solved on a computer equipped with an 8-core CPU running at 3.4 GHz and 160 GB of RAM.
@@ 4.5 Result evaluation
4.5.1 Performance metrics
Apart from the number of fast and slow chargers installed, the following performance metrics are introduced to evaluate performance:
  • Fast charging hours (FC Hrs, per truck per day): The aggregated duration, measured in half-hour intervals, that a truck uses a fast charger per day.
  • Slow charging hours (SC Hrs, per truck per day): The aggregated duration, measured in half-hour intervals, that a truck uses a slow charger per day.
  • Abandonment hours (Ab Hrs, per truck per day): The aggregated duration, measured in half-hour intervals, that a truck is abandoned and charged per day.
  • Average charging power (Avg. CP, kW): The mean power charged during all charging time slots.
  • Percentage of the number of periods when battery SoC is lower than 30% relative to the total time slots (Pct. of Low SoC, %): The total number of half-hour periods when the battery SoC is lower than 30%.
  • Average battery SoC of each truck (Avg. SoC, %): The average battery SoC of each truck throughout the trip.
  • Charging utilization rate (CUR, %): The average number of charging duration, measured in half-hour intervals, each charger experiences compared to the optimization horizon.
@@ 4.5.2 The impacts of abandonment, range anxiety, and effective parking duration
Using 39 representative days, we evaluate the impacts of abandonment, range anxiety, and effective parking duration using the HS method.
  • Case 1 - Benchmark: Trucks are prohibited from waiting in any zone under normal circumstances. However, when a vehicle’s SoC drops to below 30%, it is permitted to pause for 30 minutes to recharge. Two designated zones (i.e., zones 1 and 4) accommodate the overnight charging with unlimited waiting time. The model incorporates the specific parking duration for each parking window (i.e., $\frac{5}{30}\leq pp\leq 1$).
  • Case 2 - Full parking time: This case maintains the same as the benchmark case, except that it has no specific parking duration constraints for each parking window (i.e., $pp=1$).
@@
  • Case 4 - No range anxiety: This case maintains the same conditions as the benchmark case, except that it has no related constraints or penalty related to the range anxiety, in order words, there is no penalty for low SoC (i.e., $P^{low}=0$). Trucks are prohibited from waiting in any zone under normal circumstances, regardless of their SoC; in other words, the maximum waiting time is set to zero (i.e., $T^{max}_{i,d,t}=0$).
Figure 10: Zone-level differences in the numbers of slow and fast chargers for Cases 2–4 relative to the benchmark. Dashed bars represent fast chargers, while solid bars represent slow chargers.
Table 3 summarizes optimization results in four deterministic cases under different assumptions about abandonment, range anxiety, and effective parking duration. The benchmark case (case 1) has a mixed type of chargers, which are 72 of slow chargers and 6 of fast chargers. Fig. 10 illustrates differences in the numbers of fast and slow chargers in cases 2–4, compared to the benchmark. In the benchmark case, the number of installed chargers is proportional to the total duration of vehicle stops at each of the eight charging zones, as presented in Fig. 8 (a), while zones 4 and 5 have the highest number of chargers among the eight zones. We then examine three cases in which overnight charging, range anxiety, and limited parking duration are not considered. Extending the parking duration to the full window reduces the number of fast chargers to three. Removing the designated overnight zones would require an additional 20 slow chargers to be installed in zones 1 and 4. In case 4, no range anxiety requires seven more slow chargers to be installed.
Cases Names # of slow / fast chargers Installation costs (\$) Scheduling penalty (\$) Gap (%) Time (Hrs)
1 Benchmark 72/6 336,000 38,468 0.084 2.98
2 Full parking time 74/3 187,000 39,622 0.116 57.32
3 No overnight charging 92/6 366,000 85,200 0.042 15.77
4 No range anxiety 79/6 346,500 38,077 0.079 2.94
Table 3: Summary of charging planning outcomes and overall scheduling penalties in four deterministic cases: Benchmark, Full parking time, No overnight charging, and No range anxiety
Cases FC Hrs (truck $\cdot$ day) FCUR (%) SC Hrs (truck $\cdot$ day) SCUR (%) Ab Hrs (truck $\cdot$ day) Pct. low SoC (%) Avg. CP (kW)
1 0.18 20.64 0.50 5.46 3.60 6.01 16.51
2 0.08 19.78 0.79 8.34 3.60 4.91 13.11
3 0.19 21.65 0.48 4.06 9.46 2.12 16.91
4 0.17 19.06 0.43 4.21 3.68 62.57 19.03
Table 4: Summary of all the performance metrics in four deterministic cases; The number in bold highlights the highest values among four cases
Figure 11: The total number of (a) charging hours and (b) abandonment hours of the truck fleet in different zones in four deterministic cases
@@
Table 4 summarizes all the performance metrics in four deterministic cases, and the detailed scheduling solution of four cases are presented in Appendix 10. As expected, case 2, which assumes the full parking duration, has the highest fast-charging hours and utilization rate among the four cases. Case 3, which excludes overnight charging, has the highest number of fast-charging and abandonment hours. Case 4, with no range anxiety, leads to the highest percentage of low SoC events and the average charging power.
@@
Figs. 11 (a) and (b) show the zone-wise comparison of charging and abandonment time, respectively. In case 2, with the full parking duration, zones 6 and 8 have longer charging times. In case 3, without the overnight charging, zones 1 and 4 have a significantly high number of abandonment hours. Fig. 12 (a) illustrates the average SoC of all trucks over 39 days, ranked from the highest to the lowest, and Fig. 12 (b) presents the overall battery SoC (i.e., the maximum, minimum, and average) of the entire fleet in four cases. Among four cases, case 4 (no range anxiety) yields the lowest average SoC ($28.4\%$), whereas case 2 with full parking duration achieves the highest average SoC ($45.3\%$). When the effective charging time constraint is applied, in other words, comparing the benchmark to case 2, part of the fleet in the benchmark case experiences a sharp drop in the average SoC to very low levels ($\leq 20\%$), due to the limited charging duration in the parking windows.
4.5.3 The impacts of different RO-DDU formulations for parking duration uncertainty
We next evaluate the performance of two RO-DDU formulations for the charging power constraint under parking duration uncertainty as in Section 2.7.2. These formulations differ in whether the uncertainty set depends affinely on instantaneous charging decisions (hereafter referred to as ‘Instantaneous’) or on cluster-level expected charging decisions (hereafter referred to as ‘Expect’). We adopt fleet groups obtained from hierarchical clustering that incorporates truck category information, as described in Section 4.3. We then apply different scaling factors for slow chargers ($k_{\text{slow}}=4$) and fast chargers ($k_{\text{fast}}=2$) in the instantaneous and cluster-level RO-DDU formulations. Here, we envisage a pessimistic scenario in which parking duration will be reduced after BEV chargers are installed. In addition, we include an RO independent of charging decisions (‘RO-DI’) as a benchmark. All three cases are evaluated using 39 selected days, as defined by the IPS method. After obtaining the charger planning results, we use the sample average approximation (SAA) to compare the scheduling performance of the Instantaneous and Expect cases.
Cases Formulations # of HCVs # of slow / fast chargers Installation costs (\$) Scheduling penalty (\$) Gap (%) Time (Hrs)
1 RO - DI 19 67 / 6 328,500 20,911 0.01 0.48
2 Instant. 19 67 / 6 328,500 30,400 0.01 0.52
3 Expect 19 70 / 6 333,000 31,306 0.01 10.21
Table 5: Summary of charging planning outcomes and scheduling penalties under three RO formulations (low-penalty case)